OpenStudy (anonymous):

IS THE FOLLOWING PARALLEL, PERPENDICULAR, OR NEITHER? -3Y = -4X + 4 -4Y = 3X + 5

OpenStudy (anonymous):

Wow!!!! thank you.

OpenStudy (texaschic101):

-3y = -4x + 4 y = 4/3x - 4/3 slope = 4/3 -4y = 3x + 5 y = -3/4x - 5/4 slope = -3/4 Parallel lines will have the same slope. Perpendicular lines will have negative reciprocal slopes...all that means is " flip " the slope and change the sign. 4/3...flip and change sign = -3/4....these lines are perpendicular